What is the resistance of a 25.0-m-long piece of 12-gauge copper wire having a 2.053-mm diameter?

1
Expert's answer
2022-03-11T08:29:23-0500

The resistance of a wire

"R=\rho\frac{l}{A}\\
=1.68*10^{-8}*\frac{25.0}{3.14*(2.053*10^{-3}/2)^2}\\
=0.13\:\Omega"